Find the heat of vaporization in kJ mol-1 for compound X. The compound is found to have a vapor pressure of 15.47 kPa at 10.0 °C and 22.25 kPa at 18.0 °C.
1
Expert's answer
2020-09-17T07:14:10-0400

We know

ln( p2/p1)="\Delta\Eta" /R(T2-T1/T1T2)

Where p1=vapour pressure at T1

p2=vapour pressure at T2,

"\Delta" H= heat of vaporisation

R= universal gas constant=8.314 jule/mole k,

T1=273+10=283K

T2=273+18=291K,p1=15.47kpa,p2=22.25kpa,

Putting all these values in the above equation we can calculate the value of "\Delta" H=24.88 kj/mole